LUCIFER/LBC Technical Observing, prep for Partner Science:
UT 2010 Oct 31

Observer: DThompson, OKuhn in Tucson through 2:30am
Telescope Operator: GBechetti, SAllanson
Instrument Support: DOfficer
Telescope Support: JUrban, MMidkiff
SW support: SW Build #AO5

Summary

Preparations

Observations completed during the TCS checkout on UT 2010 Oct 30:
  • Works under BP5 TCS
  • UI was showing updates of data from TCS
  • Able to preset and offset from the UI with no issues

Details

Reconfigured for LBC to check collimation if weather allows (currently windy)

Data: luci.20101031.NNNN.fits

rangebal task in Observe package tested in closed dome.
  • works for L/R data based on LBTtools.Telside
  • Corrected M2 RX/RY mis-calculation
  • test on sky ASAP

02:18 LUCIFER darks (use LUCIFER$/readnoise script)
  • O2DCR 2.0s (0001-0002) 11.42 e-
  • MER10 10.0s (0003-0004) 5.41 e-

Wrote slitdither.*, imagepair, and dice5_10x generic scripts (tested below)

06:30 Opened dome. Winds going down slowly. T=4.2, RH=21 mostly clear.

ACT0366
IDL> dofpia, /first (red nearly in focus, dofpia failed to collimate. -0.5mm Z global applied to red side)
IDL> dofpia, /first, /backout
IDL> dofpia, /first (hit limits on RED side)

Hit limits a couple times on RED and had to add another +1000 z11. We added a global offset of -0.5mm Z to DX and the total collimation Z=-2.5mm so we will likely need to crank the red hub up somewhat to avoid hitting that limit. Collimation reasonably close. "S" to stop.

07:15 WT10_327 (EL~84)
IDL> dofpia
Immediately hit a Z limit on RED. The hub will have to be moved. Cleared all active optics and forces.
IDL> dofpia, /redonly, /first

RED hub will have to be moved. We (OKuhn, DThompson) recommend +2 mm to get to parity with where BLUE was tonight. DOfficer's or JHill's memory can modify this. Email will be sent to telescopework, DOfficer, MWagner with this request. Given that it should get quite a bit colder than tonight's 4C, we may want to add +1mm to both on top of this (+1mm BLUE, +3mm RED).

07:41 Switching to LUCIFER+IRTC

08:28 Pupil imaging on sky, EL~81 degrees. Files 0007,08 = PARANG 0 degrees, 0009,10 = 90d, 0011,12 = 180d, 0013,14 = 270d, 0011 = 0 (NDIT=2, normal). Flexure compensation was off, some of the pupil motion could be because of that. Jpegs sent to MWagner, JHill, WSeifert, NAgeorges 31 Oct 18:00 MST.

08:39 Correct initial collimation/pointing ACT0342 to correct pointing. File 0017. IE changed from -81 to -79.5, CA from 20. to 23.6. Initial collimation looks decent (stars are round and not obviously out of focus or lots of coma)
BS9107 to collimate. 1" FWHM variable.

Astrometric data BS9107 Files 0018, 0019 BrG+Ks 5x10.0s integ.

Check existing collimation model BS9107 EL~80 collimated
BS9122 EL~56 looks OK, pointing off
BS9126 EL~50 looks OK
BS9131 EL~40 starting to be bad but still compact, no problem for AGw

Check of dithering along slit (10x 2' dither) Preset to available BS9122 star from UI/Catalog
09:13 Execute script: ~/LBTO/SciencePrep/slitdither.acq
Guide star fading in and out from clouds!
Align BS source to center of 2" longslit (files 0022-30)
Execute script: ~/LBTO/SciencePrep/slitdither.sci (Files 0031-50)
Note: FC not on per LUCI Team request
Source stayed well-centered in slit. Quantitative results to follow

Astrometric field data to check AGw calibration for monitoring
Execute script ~/LBTO/SciencePrep/imagepair (Files 0051-0052) BS9122 5x10;0s Ks

Checking wind speed by rotating the enclosure

Take a set of (monocular) range-balanced collimation data Windy, clouds passing overhead, seeing not great.

BS9122 EL~69 allow to collimate, then rangebal then take image to record collim data

BS9122 EL~69 0053
BS9131 EL~46 0054
BS9139 EL~48 0055? (no luci image taken, get mirror pos from notes)
BS9122 EL~73 00xx
BS9126 EL~77 0056
BS9139 EL~50 0057
BS9142 EL~50 0058

Take a set of (monocular) range-balanced collimation data
Windy, clouds passing overhead, seeing not great.

BS9142 EL~53 0059
AO378 EL~45 0060
AO379 EL~37 0061
BS9142 EL~56 0062
AO320 EL~64 0063
AO330 EL~71 0064
AO306 EL~80 0065
AO304 EL~83 0066
AO309 EL~76 0067
AO309 EL~76 0068?
AO366 EL~66 0069
AO378 EL~59 0070

13:03 lost guiding because of twilight. Closing dome.

Observing Plan

  • Twilight flats in NB, BB filters (cycle through all over time)
    • Need to set up a records page on the wiki to keep track

  • Test any new template observing scripts prior to release

  • Repeat measurement of focal plane scale, orientation

  • Additional work on Andrew's M1/M2 collimation?

  • Check existing pointing model
    • Implement new collimation model first, if necessary
    • Stable conditions!

  • Take a set of pointing data using range-balanced collimation model
    • Stable conditions!

  • Collect field aberration data

  • Photometric zero point and extinction monitoring

  • Collect a "movie" of 2s LUCIFER images while an in-field offset is taking place.
    • Check to make sure offsets still happen in a reasonable manner

  • Collect through-focus sequence for Andrew - reasonably dense star field.
    • Try Z4 of -1000, -500. 0, 500, 1000 nm
    • Also check the AGw-LUCIFER focus in cold weather for temp term?

  • Collect astrometric field data to check AGw calibration for monitoring

  • Improve map of AGw patrol field limits at high field angle
    • Requires set up work on a stone field

  • Gather NB illumination correction data (BrG done on 20100606)

  • Re-test updatePointingReference functionality.

  • Continue gaining experience with LUCIFER on-sky (e.g. spec acq.)

  • Test dithering between extreme slits on MOS mask + telluric
    • Manual needs 0.59836mm/" scale for the dXdY calculation
    • The how-to set up telluric observations needs updating for clarity.

  • Investigate variable imaging FF gradients by taking data while the telescope is offset with respect to the enclosure (will need manual overriding of normal operations) to see if the observed gradients could be caused by something 'external to the instrument'.

  • Roy van Boekel's test observations of wasp12b

-- DavidThompson - 22 Oct 2010
Topic revision: r6 - 01 Nov 2010, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback